OpenStudy (anonymous):

what is y=1/2x+3 written standard form using integers?

OpenStudy (therealmeeeee):

2x + 5y = –15 2x + y = –3 2x + 5y = –3 x + y = –15

OpenStudy (therealmeeeee):

This just depends on what country your from! I think you want this: y=3x+1 3x - y = -1 ... I've seen people do this and this website calls it standard form http://www.algebralab.org/studyaids/stud... or do you want 3x - y + 1 = 0?? .. we call this "general form" y=1/2x-3 2y = x - 3 x - 2y = 3 y=-3x-4 3x + y = -4 y=7/2x+1/4 4y = 14x + 1 14x - 4y = -1 3y=-3x y = -x x + y = 0

OpenStudy (therealmeeeee):

do this help @precious13

OpenStudy (anonymous):

ok thanks @TheRealMeeeee

OpenStudy (therealmeeeee):

click best response

OpenStudy (therealmeeeee):

r u going to click it

OpenStudy (therealmeeeee):

ty

OpenStudy (therealmeeeee):

need my help let me know umm tag me